Find all solutions of the equation x^2−3x+8=0 and express them in the form a+bi:
First with the input b<0
Then with the input b>0

Answer by stanbon(75887)

x^2−3x+8=0
-----
Use the Quadratic Formula to get:
x = [3+-sqrt(9-4*8)]/2
---
x = [3+-sqrt(-23)]/2
----
x = (3/2) + (sqrt(23)/2)i or x = (3/2)-(sqrt(23)/2)i
======================================================
Cheers,
Stan H.
========